OpenStudy (anonymous):

A line contains points (-4, 2) and (0, -4). Another line graphed in the same coordinate plane contains points (4, 3) and (-5, -3). Based on the slopes of these lines, are they parallel, perpendicular or neither?

OpenStudy (anonymous):

Parallel lines: same slope Perpendicular lines: negative inverse slope

OpenStudy (anonymous):

inverse slope

OpenStudy (anonymous):

its Parallel

OpenStudy (anonymous):

What did you get for the slope of the first line?

OpenStudy (anonymous):

Inverse slope means they are perpendicular. For Example If a line has slope 3 and other has -1/3 then these two lines are perpendicular.

OpenStudy (anonymous):

so it will be Parallel

OpenStudy (anonymous):

What did you get for the slopes?

OpenStudy (anonymous):

Yes if the slopes of both the lines are same then lines are parallel.

OpenStudy (anonymous):

??????????????

OpenStudy (anonymous):

You have to find the slope of both lines, do you know how to do that?

OpenStudy (anonymous):

no it has to be perpendicular

OpenStudy (anonymous):

2/4

OpenStudy (anonymous):

I gotten 2/4 and 1

OpenStudy (anonymous):

-3/2 and 2/3. 2/3 is negative reciprocal of -3/2.So Lines are perpendicular.

OpenStudy (anonymous):

how 2/4 and 1 lvasquev462? when there is a formula to find slope which is: y2-y1/x2-x1. So -4-2/0-(-4)=-6/4=-3/2

OpenStudy (anonymous):

Similarly find the slope of second equation.Which I got as 2/3. So (m1)(m2)=-1.Then lines are perpendicular. where m1=slope of first equation and m2=slope of second equation.

OpenStudy (anonymous):

You remember formula (m1)(m2)=-1 to check either lines are perpendicular or not.ok